|x-1|+|x-2|+|x-3|=2.what is the solution of the equation

Grade:11

1 Answers

Nidhi
6 years ago
x=8/3 is solution of equation..by adding like terms we get 3x-6=2, 3x=2+6,x=8/3.. So this is the answer..